OpenStudy (anonymous):

Solve each simultaneous equation X+Y=264 X-Y=208

OpenStudy (anonymous):

add both of them to get x, then substitute in any of the eqns to get y....

OpenStudy (anonymous):

So add the two numbers together

OpenStudy (anonymous):

nope add the two equations both the sides, y gets cancelled then simplify to get x

OpenStudy (anonymous):

Can you show me??

OpenStudy (anonymous):

ahhhhhh... x+y=264 x-y=208 2x=264+208 2x=272 x=136 put it in y y=264-x y=264-136 y=128 hope its correct

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

2x=264+208 2x=472 x=236 when u substitute in y=264-x y=264-236 =38